5) A certain standardized test has scores which range from 0 to 500, with decimal scores possible. Scores on the exam are normally distributed with a mean of 319 and a standard deviation of 46.

What proportion of students taking the exam receive a score greater than 368?

Round your answer to 4 decimal places.
Answer #1

Solution :

Given that ,

mean = = 319

standard deviation = = 46

P(x > 368) = 1 - P(x < 368)

= 1 - P[(x - ) / < (368 - 319) / 46)

= 1 - P(z < 1.0652)

= 1 - 0.8566

= 0.1434

proportion = 0.1434